# Attenuation of the ultra-high-energy neutrino flux by dark matter scatterings

### v1

@@ 5 Conclusions
We have presented a comprehensive analysis of the attenuation of a diffuse neutrino flux due to scattering with DM particles in the intergalactic medium and in the Milky Way halo. We have argued that when the scattering cross-section over DM mass is sizable, the attenuation in the Milky Way generates a neutrino flux with a characteristic anisotropy.
We have then studied the implications of DM-$\nu$ interactions for the flux, energy spectrum and arrival direction of ultra-high energy neutrinos, assuming that they are of diffuse origin. Concretely, we have shown that the Waxman-Bahcall limit becomes more stringent in the presence of DM-$\nu$ interactions. Further, we have investigated the implications of the detection of the event KM3-230213A for the DM-$\nu$ interactions, assuming that the event is of diffuse origin. If this is the case, future instruments should detect hundreds of ultra-high-energy neutrinos, which will allow to perform a detailed spectral and morphological analysis of the signal, which encode invaluable information about the magnitude and energy dependence of the scattering cross-section. Finally, we have highlighted the complementarity of different neutrino detectors at different latitudes in the Earth, in order to reveal the existence of particle DM through the anisotropy in the arrival directions of ultra-high-energy neutrinos due to their passage through the DM halo of our Galaxy and the Earth.
@@ Appendix A Evaluation of the effective areas
We compute the angular dependence of the IceCube and KM3NeT effective areas following Ref. [37]. The effective area for a neutrino with energy $E_{\nu}$ and zenith angle $\theta$ that interacts inside the detector is given by
@@
with $\sigma$ the neutrino-nucleon interaction cross section, $N_{\mathrm{nuc}}$ the number of nucleons in the detector, $\varepsilon$ the detection efficiency, and $n(s,\theta)$ the number density of nucleons along the traversed Earth path. The last term factor into account neutrino attenuation by Earth. In all our calculations, we assume the Preliminary Earth Reference Model (PREM) [38].

### v2

@@ 5 Conclusions
We have presented a comprehensive analysis of the attenuation of a diffuse neutrino flux due to scattering with DM particles in the intergalactic medium and in the Milky Way halo. We have argued that when the scattering cross section over DM mass is sizable, the attenuation in the Milky Way generates a suppressed neutrino flux with a characteristic anisotropy.
We have then studied the implications of DM-$\nu$ interactions for the flux, energy spectrum and arrival direction of ultra-high energy neutrinos, assuming that they are of diffuse origin. Concretely, we have shown that the Waxman-Bahcall limit becomes more stringent in the presence of DM-$\nu$ interactions. This implies that, with mild astrophysical assumptions, limits on the DM-$\nu$ interaction cross section can be placed even if the unattenuated flux is unknown. Further, we have investigated the implications of the detection of the event KM3‑230213A for the DM-$\nu$ interactions, assuming that the event is of diffuse origin. This allows to exclude DM-neutrino interaction cross sections $\sigma/M_{\mathrm{DM}}\gtrsim 4\times 10^{-22}\,\mathrm{cm^{2}/GeV}$ at unprecedentedly high energies. If the event is indeed of diffuse origin, future instruments should detect hundreds of UHE neutrinos, which will allow to perform a detailed spectral and morphological analysis of the signal, that encode invaluable information about the magnitude and energy dependence of the scattering cross section. Finally, we have highlighted the complementarity of different neutrino detectors at different latitudes in the Earth, in order to probe DM-neutrino interactions through the anisotropy in the arrival directions of UHE neutrinos due to their passage through the DM halo of our Galaxy and the Earth.
@@ Appendix A Evaluation of the effective areas
We compute the angular dependence of the IceCube and KM3NeT effective areas following Ref. [44] (see Ref. [45] for a more detailed approach using a transport equation). The effective area for a neutrino with energy $E_{\nu}$ and zenith angle $\theta$ that interacts inside the detector is given by
@@
with $\sigma$ the neutrino-nucleon interaction cross section, $N_{\mathrm{nuc}}$ the number of nucleons in the detector, $\varepsilon$ the detection efficiency, and $n(s,\theta)$ the number density of nucleons along the traversed Earth path. The last term takes into account neutrino attenuation by Earth. In all our calculations, we assume the Preliminary Earth Reference Model (PREM) [46].